CH/Agen/93/02

Agenda

Charissa Group Meeting

To be held on: Wednesday, 7 April 1993, 11.15 am To be held at: Daresbury Laboratory

Distribution: WNC,NMC,RAC,BRF,GJG,SJAH,RAH,JSL,ACM, VP,WDMR,GT,RPW,DLW,JZ

1. Minutes of Meeting of 10 February 1993
2. Matters Arising/ Actions
3. Update on NSC/SERC matters  (DLW)
4. Progress on detector tests  (DLW)
5. Progress on chamber & detector assemblies  (JSL)
6. Progress on gas supply  (RAC)
7. Progress on electronics & data acquisition  (RAH)
8. Developments on event manager & GEC emulator  (VP)
9. Requirements for hybrid detector operation at the ANU  (RAC)
10. Report on Planning Trip to the ANU  (BRF,JSL)
11. Timetabling of manpower for ANU and selection of experiments
11. Packing for shipment to the ANU  (DLW)
13. Documentation for ANU Running
14. Plans for Strasbourg
18. Any other business
19. Date of next meeting

W.N. Catford (Secretary) 24 March 1993